What is the separable variable? Dy/dx = ye^-x^2, y(4) = 1?
1
Expert's answer
2021-10-19T10:20:27-0400

Solution:

A differential equation is said to be separable variable if the variables can be separated.

Given,

"\frac{dy}{dx} = ye^{-x^2}"

"\Rightarrow \frac{dy}{y} = e^{-x^2}dx"

"\Rightarrow \int\frac{dy}{y}=\int e{^{-x^2}}dx"

But "\int e^{-x^2}dx"  is a non elementary integral. We can express it as a power series

"e^{-x^2}= 1 - \frac{x^2}{1!}+\frac{x^4}{2!}-\frac{x^6}{3!} +\frac{x^8}{4!}-.."

We can approximate it as

"e^{-x^2}= 1 - \frac{x^2}{1!}+\frac{x^4}{2!}-\frac{x^6}{3!}"

So differential equation transforms to

"\int\frac{dy}{y}=\int" "(1 - \frac{x^2}{1!}+\frac{x^4}{2!}-\frac{x^6}{3!})dx"

 "\Rightarrow \ln |y|=x - \frac{x^3}{3}+\frac{x^5}{10}-\frac{x^7}{42} + C"

By initial condition y(4)=1

0 = "4 - \frac{4^3}{3}+\frac{4^5}{10}-\frac{4^7}{42} + C"

"\Rightarrow C=305.028"

So ln|y| = "x - \frac{x^3}{3}+\frac{x^5}{10}-\frac{x^7}{42} +305.028"

 "\Rightarrow y=e^{(x - \frac{x^3}{3}+\frac{x^5}{10}-\frac{x^7}{42} + 305.028)}"

This is the explicit solution of the given differential equation.